OpenStudy (anonymous):

520 + 40x = 3800 - 120y

hero (hero):

Solve for which variable?

OpenStudy (anonymous):

x

hero (hero):

520 + 40x = 3800 - 120y 40x = 3800 - 520 - 120y 40x = 3280 - 120y x = (3280 - 120y)/40

OpenStudy (anonymous):

so what is x?

OpenStudy (anonymous):

like when i plug it in, it's going to work

OpenStudy (anonymous):

x = 82 - 3 y

OpenStudy (anonymous):

what is it exactly? lol

OpenStudy (anonymous):

Simplifying 520 + 40x = 3800 + -120y Solving 520 + 40x = 3800 + -120y Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-520' to each side of the equation. 520 + -520 + 40x = 3800 + -520 + -120y Combine like terms: 520 + -520 = 0 0 + 40x = 3800 + -520 + -120y 40x = 3800 + -520 + -120y Combine like terms: 3800 + -520 = 3280 40x = 3280 + -120y Divide each side by '40'. x = 82 + -3y Simplifying x = 82 + -3y so plug x in Simplifying 520 + 40(82 + -3y) = 3800 + -120y 520 + (82 * 40 + -3y * 40) = 3800 + -120y 520 + (3280 + -120y) = 3800 + -120y Combine like terms: 520 + 3280 = 3800 3800 + -120y = 3800 + -120y Add '-3800' to each side of the equation. 3800 + -3800 + -120y = 3800 + -3800 + -120y Combine like terms: 3800 + -3800 = 0 0 + -120y = 3800 + -3800 + -120y -120y = 3800 + -3800 + -120y Combine like terms: 3800 + -3800 = 0 -120y = 0 + -120y -120y = -120y Add '120y' to each side of the equation. -120y + 120y = -120y + 120y Combine like terms: -120y + 120y = 0 0 = -120y + 120y Combine like terms: -120y + 120y = 0 0 = 0 Solving 0 = 0 This equation is an identity, all real numbers are solutions.